OpenStudy (anonymous):

alex opened a new savings account and she planned to save $550 each month. She was however, able to save more each month than she had originally planned. In fact after 12 months (without considering interest) Alex's savings account will be three months short of twice the amount she originally thought she would save in 12 months. How much did Alex actually save each month?

OpenStudy (saifoo.khan):

1. if she is 3 months short, then in 15 months, she will have saved twice the amount she originally thought. Algebraically, this means that if x is the actual amount she put away, 15x = 2(12*550) the RHS means take 550 per month for one year and double it, which is what the sentence says. Solve for x and there ya go.

OpenStudy (anonymous):

$550(12)=6600 2(6600)=13200 13200-(3)550=11550 11550/12= $962.50